# Towards Better Open-Ended Text Generation: A Multicriteria Evaluation Framework

### v1

@@ 3.2 Evaluation Metrics
3.3 Datasets
Furthermore, we utilize the dataset provided by Garces-Arias et al. (2024a), which includes over 2.2 million generated continuations (with a maximum length of 256 tokens) for each prompt, along with aggregated metrics such as coherence, diversity, and MAUVE. We extend this dataset by computing sentence-level evaluations and incorporating generation perplexity, as described above in subsection 3.2.
3.4 Decoding strategies and Hyperparameters
4 Benchmarking utilizing order structure
In this section, we propose a benchmarking of decoding strategies along the Bradley-Terry model Bradley and Terry (1952b) and the union-free-generic (ufg) depth Blocher et al. (2024); Blocher and Schollmeyer (2024). Both approaches deliver rankings of decoding strategies rather than metric values for each strategy. We motivate this from a very practical perspective, see section 1: In many cases, the metric information can be considered redundant. That is, ranking the decoding strategies is the overall aim of benchmarking, not assigning metric values. After all, a decoding method can either be deployed or not by practitioners, rendering the metric information not of direct practical interest.
4.1 Extended Bradley-Terry model
The extended Bradley-Terry model is based on pairwise comparisons, see Bradley and Terry (1952a); Davidson (1970). To obtain this pairwise comparison of the methods, we compare the methods pairwise on each single generation. For simplicity, let us look at WikiText-103 with 1314 generations. For each generation $G_{i,j}$ (from model $i$ with decoding strategy $j$, see Figure 1), we can evaluate the methods based on coherence, perplexity, and diversity. Now we consider each pair of methods and each generation, e.g. Mistral 3 CS ((’0.2’, ’1’)) and Mistral 3 CS ((’0.8’, ’1’)) with generation id 1. Then we state that method 1 dominates method 2 if and only if there exists one metric that strictly prefers method 1 over method 2 and the other two metrics do not disagree. Based on generation id 1, we obtain that Mistral 3 CS ((’0.2’, ’1’)) is preferred over Mistral 3 CS ((’0.8’, ’1’)) since coherence states that method Mistral 3 CS ((’0.2’, ’1’)) is preferred over method Mistral 3 CS ((’0.8’, ’1’)), and the other two metrics, here diversity and perplexity, do not contradict this statement. In this way, for WikiText-103 we get 1314 comparison statements for each pair of two methods. Note that when two methods have exactly the same metric values, they are considered indifferent. This differs from incomparability; for details, see Rodemann and Blocher (2024). However, with slight abuse, we do not distinguish these two cases here.
4.2 Union-Free Generic Depth
To account for exactly this complexity, we build on Rodemann and Blocher (2024); Blocher et al. (2024) and deliberately allow for incomparability. Mathematically, this means that we aim for partial rather than total orders. Let us look again at a single generation $G_{i,j}$. For the extended Bradley-Terry model, we only considered the pairwise comparisons. However, all the pairwise comparisons resulting from a generation $G_{i,j}$ together define a partial order that describes the preference structure of the methods based on that generation $G_{i,j}$. This gives us 1314 partial orders for the WikiText-103 data. For example, two observation are now the entire partial orders in Figure 2.
5 Benchmarking by collapsing into a single metric
As previously discussed, metrics such as coherence and diversity, when optimized excessively, can lead to unintended outcomes. High coherence scores often result in repetitive or degenerate text, as the model overemphasizes continuity at the expense of introducing new information. Conversely, prioritizing high diversity can produce erratic sequences that lose relevance to the initial prompt, undermining the overall quality of the generated text. The construction of Q*Text involves three key steps: normalization to a common scale of 0 to 100, where 0 represents the worst performance and 100 represents the best, penalization of extreme values, and aggregation using the harmonic mean. This section details each step and the rationale behind our methodological choices.
5.1 Normalization
@@ 5.3 Aggregation
6 Results
The extended Bradley-Terry model returns so-called "worth" parameters which indicate the probability that this decoding method is preferred over the other in a pairwise comparison. When all datasets are considered at once, the method that dominates all other methods according to the extended Bradley-Terry model is Mistral 3 CS ((’0.4’, ’10’)). The second best method is Mistral 3 CS ((’0.4’, ’5’)). The worst method is GPT2-XL CS ((’1.0’, ’20’)). A snippet of the results, including the case when restricting the analysis to only one dataset, can be found in Appendix A.2.
Therefore, in the next step, we consider the union-free generic depth approach, which allows for two methods to be incomparable. Furthermore, the ufg-depth considers the entire set of pairwise comparisons for a generation as one observation and does not assume an independence structure between them. Due to the high computational complexity, see Appendix A.3, we restrict our analysis to the WikiText-103 dataset and compare only the four methods that appear to be the best according to the extended Bradley-Terry model, see Appendix A.2: Mistral 3 CS ((’0.6’, ’15’)), Mistral 3 CS ((’0.4’, ’3’)), Mistral 3 CS ((’0.8’, ’3’)), Mistral 3 CS ((’0.4’, ’20’)) and Mistral 3 CS ((’0.4’, ’50’)).
The highest ufg-depth with a value of 0.977 (thus the one that has the structure most supported by the observation), is the one that shows no dominance structure among the four methods, i.e. the one that concludes that all methods are incomparable to each other, see Figure 2 (top). The partial order with the lowest depth value of 0.003 prefers Mistral 3 CS ((’0.4’, ’3’)) over Mistral 3 CS ((’0.4’, ’20’)), see Figure 2 (bottom). This order can be considered an outlier. In other words, it is the partial order that has the least support by the benchmarking results.
With regards to Q*Text, we observe that balanced deterministic strategies such as contrastive search with moderate penalties ($\alpha$ values of 0.4 or 0.6) and moderate values of $k$ (5, 10, or 15) are preferred, as illustrated in Tables 16 and 14. Combinations that counterbalance each other—for instance, low $\alpha$ values (0.2) with high $k$ values (20 or 50), or high $\alpha$ values (0.8 or 1.0) with moderate $k$ values (3 or 5)—also yield favorable results. Both Q*Text and the extended Bradley-Terry model indicate a general preference for larger architectures; however, smaller models like GPT2-XL can outperform more modern architectures when paired with balanced decoding strategies, see Table 10.
To evaluate the level of agreement between the total ordering from the extended Bradley-Terry and that provided by Q*Text, we analyze and visualize the decoding methods with the highest agreement and those with the largest rank deviations (see Appendix A.4, Figures 3 and 4). The results suggest discrepancies, particularly for less diverse and less coherent generations. Nevertheless, there is good agreement when ranking decoding methods with moderate hyperparameters, as previously discussed.
Despite operating on different underlying mechanisms, the preferences from the extended Bradley-Terry model exhibit noteworthy similarities with those from Q*Text concerning the aforementioned balanced strategies. One notable difference is that the extended Bradley-Terry model does not account for significant drops in diversity, as seen with beam search, which reduces the penalization of this decoding strategy. On the other hand, decoding methods such as GPT2-XL with contrastive search parameters ($\alpha$ of 1.0 and $k$ of 20) are strongly penalized by both approaches due to their incoherence and low generation confidence, see Tables 11, 13 and 17.
Finally, we naturally ask how these alternative methods, evaluated within a multicriteria framework, align with human preferences. To address this, we considered the rankings provided by human evaluators in Section 5 of Garces-Arias et al. (2024a), which present similar findings. Specifically, human evaluators favored balanced contrastive search strategies, e.g., (0.4, 10), (0.4, 20), (0.4, 5), (0.6, 10), and (0.6, 5), top-$k$ sampling with $k=10$ or 50, and nucleus (top-$p$) sampling with $p=0.9$. In contrast, unbalanced contrastive search approaches, such as (1.0, 50) and (0.8, 50), were consistently evaluated as low quality due to their low coherence. Beam search strategies with moderate beam widths, e.g. less than or equal to 20, were also often judged as low quality due to their low diversity. See Tables 8, 9 and 16 and the work by Garces-Arias et al. (2024a) for more details. Furthermore, Table 18 presents examples of generations from various decoding methods, highlighting those that received positive and negative evaluations from Bradley-Terry and Q*Text.
7 Conclusion
By applying these methodologies in extensive experiments across diverse datasets and models, we have shown that our proposed approach not only enhances the interpretability of text generation evaluations but also provides a practical, data-driven framework for benchmarking algorithms in this domain. We hope that these insights will contribute to a more rigorous evaluation of text generation models and inspire further research in the field of multicriteria benchmarking strategies.
Key Takeaways and Recommendations.
Summing things up, our study revealed that different practical scenarios require different multicriteria benchmark evaluation frameworks. Instead of relying on one single benchmark suite with a pre-specified evaluation method, we recommend practitioners to define the overall aim of benchmarking and evaluation thereof as precisely as possible. Specifically, we identify two crucial questions to be answered beforehand:
  1. 1.
    Is it sufficient to rank methods, or is metric information about the methods’ performances required?
  2. 2.
    Does the use case require a total or partial ordering method, i.e., should the evaluation allow for incomparability among some methods, or should it enforce comparability of all methods?
In case metric information is required and comparability of all methods should be enforced, we recommend our novel aggregation metric Q*Text, see section 5. If the metric information is not the overall aim, but comparability should still be enforced, we recommend using the Bradley-Terry model, see subsection 4.1. Eventually, if a ranking is required that allows for incomparability, we recommend deploying ufg-depth; see subsection 4.2.
Limitations
Ethics Statement
Acknowledgments
Matthias Aßenmacher received funding from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) as part of BERD@NFDI, under grant number 460037581. Julian Rodemann acknowledges support by the Federal Statistical Office of Germany within the co-operation project "Machine Learning in Official Statistics" as well as by the Bavarian Institute for Digital Transformation (bidt) and the Bavarian Academy of Sciences (BAS) within a graduate scholarship.
Appendix A Appendix
A.1 Result of pairwise comparison
@@
When we merge the three datasets WikiText-103, Wikinews and Book, we consider $1314+2000+1947=5261$ generations and 124962 pairwise comparisons based on each generation. Comparing the tables 2, 3, 4 we find that there is no pairwise comparison that occurs in each table. Therefore, there is no pair of two methods where method 1 dominates method 2 based on all 5261 generations. With 4601 is the dominance of Mistral 3 CS ((’0.8’, ’10’)) over GPT2-XL CS ((’1.0’, ’10’)) the one that occurs most often. For 2990 pairwise comparison at least on 90% of the generations method 1 dominates method 2 strictly. In 9191 pairwise method comparisons, we obtain that method 1 never strictly dominates method 2 (i.e. on every generation method 2 either dominates method 1 or the three metrics disagree on the dominance structure or a completely equal).
A.2 Results of the extended Bradley-Terry model
@@
For reasons of clarity and comprehensibility, we decided to show here only a snippet, but the full result can be easily and fast obtained by the already stored results in GitHub-repository. Table 5 denotes the worth parameter based on WikiText-103, Table 6 on Wikinews, Table 7 on Books and all three datasets combined can be seen in Tables 8 and 9. All computations are based on the metrics of perplexity, coherence, and diversity.
@@
A.3 Computation Complexity of the Algorithms
A.4 Results of Q*Text
Based on the Q*Text metric introduced in Section 5, we can induce a total ordering of decoding methods. Tables 10, 12, 14 and 16 illustrate the results for the most dominant decoding models, strategies, hyperparameters and methods, respectively. On the other hand, We observe in Tables 11, 13, 15 and 17 the results for the least dominant decoding models, strategies, hyperparameters and methods.
@@ A.4 Results of Q*Text
Prompt "Headlam served as Officer Commanding North @-@ Western Area in 1946, and as Director of Training from 1947 to 1950. In 1950 - 51,",
@@
Mistral 3 CS (0.4, 10) "he was Chief of Staff of the Malayan Armed Forces. Headlam was promoted to Major General in 1951, and served as GOC Far East Land Forces from 1951 to 1953. He was promoted to Lieutenant General in 1953, and served as Adjutant General from 1953 to 1956…. (text continues)"
@@
Mistral 3 CS (0.4, 5) "he was Commandant of the Royal Military Academy Sandhurst. He was promoted to Major General in 1951, and served as General Officer Commanding 1st Division from 1951 to 1953. Headlam was appointed Chief of the Imperial General Staff in 1953, and held that post until 1956. He was promoted to Field Marshal in 1956, and retired from the Army in 1957 … (text continues)"
@@
Deepseek CS (0.6, 15) "he was Chief of Staff (C.O.S.) of the 1st Parachute Brigade, and in 1951 2̆013 52, he was G.O.C. of the 1st Parachute Brigade. He was promoted to Brigadier… (text continues)"
@@
Falcon 2 CS (0.8, 1) "he was the first Commandant of the Staff College, Camberley. He was appointed General Officer Commanding (GOC) 1st Division in 1951, and GOC 3rd Division in 1953. He was appointed Commander-in-Chief (C-in-C) of the British Army of the Rhine (BAOR) in 1955, and retired in 1958… (text continues)"
@@
GPT2-XL CS (0.6, 5) Headlam served as Assistant Chief of Staff, Headquarters U.S. Army (AHCUS) at Fort Leavenworth, Kansas, before being promoted to Brigadier General and assigned to the United States Army Air Forces (USAAF) at Ramstein Air Base, Germany. After the Korean War, Headlam was a consultant to the U.S. Strategic Air Command (SAC) and the National Reconnaissance Office (NRO, which was established in the 1950s to develop intelligence satellites and spy planes).
@@
GPT2-XL CS (1.0, 20) "Headless Horsemurdertrud (his nom de v0̆0eal et re, a translators and transcriph.coastavarism in Latin-Ceux du Ponto-Monahawak at Keremej, New Heythorhampsthembryghtes-sur Meriont Bay of Sulaunkhuyut/Arap Moheluwempeh is one hundred nautilus middelveth the seagoch that swim to an anchord-like… (text continues)"
@@
Qwen 2 CS (1.0, 20) " Headelm oversa 4̆e86 Operation Agulhene535af69, to reoccupyo portug 1 st colonial outst a ncesin AFRicacoun trie sto preva l sepa rabli st mov me s . In th es on he wro 4he couthoea2 rsh ipp ublicatin hitherto re t aine o secret… (text continues)"
@@
Llama 3 beam (3) "he served as Commanding Officer of the 1st Battalion, The Queen’s Royal Regiment (West Yorkshire Regiment). In 1951 2013 52, he served as Commanding Officer of the 2nd Battalion, The Queen’s Royal Regiment (West Yorkshire Regiment). In 1952 2013 53, he served as Commanding Officer of the 3rd Battalion, The Queen’s Royal Regiment (West Yorkshire Regiment). In 1953 2013 54, he served as Commanding Officer of the 4th Battalion, The Queen’s Royal Regiment … (text continues)"
@@
Human "he was Director of Operations and Intelligence, and in 1951–54, Commander of the 1st Division, which was the most powerful division in the world. He was appointed Commander-in-Chief of the Army in 1954… (text continues)"

### v2

@@ A.1 Automatic metrics
A.2 Q*Text
Q*Texts results from the harmonic mean of normalized and penalized coherence, diversity, and generation perplexity (see Section A.1).
A.2.1 Normalization
@@ A.2.3 Aggregation
A.3 Result of pairwise comparison
@@
When we merge the three datasets WikiText-103, Wikinews and Book, we consider $1314+2000+1947=5261$ generations and 124962 pairwise comparisons based on each generation. Comparing the tables 4, 5, 6 we find that there is no pairwise comparison that occurs in each table. Therefore, there is no pair of two methods where method 1 dominates method 2 based on all 5261 generations. With 4601 is the dominance of Mistral 3 CS ((’0.8’, ’10’)) over GPT2-XL CS ((’1.0’, ’10’)) the one that occurs most often. For 2990 pairwise comparison at least on 90% of the generations method 1 dominates method 2 strictly. In 9191 pairwise method comparisons, we obtain that method 1 never strictly dominates method 2 (i.e. on every generation method 2 either dominates method 1 or the three metrics disagree on the dominance structure or a completely equal).
A.4 Results of the extended Bradley-Terry model
@@
For reasons of clarity and comprehensibility, we decided to show here only a snippet, but the full result can be easily and fast obtained by the already stored results in GitHub-repository. Table 7 denotes the worth parameter based on WikiText-103, Table 8 on Wikinews, Table 9 on Books and all three datasets combined can be seen in Tables 10 and 11. All computations are based on the metrics of perplexity, coherence, and diversity.
@@
A.5 Discussion of the Ufg-depth Results
A.6 Results of Q*Text
Based on the Q*Text metric introduced in Section 5, we can induce a total ordering of decoding methods. Tables 12, 14, 16 and 18 illustrate the results for the most dominant decoding models, strategies, hyperparameters and methods, respectively. On the other hand, We observe in Tables 13, 15, 17 and 19 the results for the least dominant decoding models, strategies, hyperparameters and methods.
@@ A.6 Results of Q*Text
Prompt "Headlam served as Officer Commanding North @-@ Western Area in 1946, and as Director of Training from 1947 to 1950. In 1950 - 51,", Q*Text
@@
Mistral 3 CS (0.4, 10) "he was Chief of Staff of the Malayan Armed Forces. Headlam was promoted to Major General in 1951, and served as GOC Far East Land Forces from 1951 to 1953. He was promoted to Lieutenant General in 1953, and served as Adjutant General from 1953 to 1956…. (text continues)" 81.62
@@
Mistral 3 CS (0.4, 5) "he was Commandant of the Royal Military Academy Sandhurst. He was promoted to Major General in 1951, and served as General Officer Commanding 1st Division from 1951 to 1953. Headlam was appointed Chief of the Imperial General Staff in 1953, and held that post until 1956. He was promoted to Field Marshal in 1956, and retired from the Army in 1957 … (text continues)" 77.63
@@
Deepseek CS (0.6, 15) "he was Chief of Staff (C.O.S.) of the 1st Parachute Brigade, and in 1951 2̆013 52, he was G.O.C. of the 1st Parachute Brigade. He was promoted to Brigadier… (text continues)" 80.16
@@
Falcon 2 CS (0.8, 1) "he was the first Commandant of the Staff College, Camberley. He was appointed General Officer Commanding (GOC) 1st Division in 1951, and GOC 3rd Division in 1953. He was appointed Commander-in-Chief (C-in-C) of the British Army of the Rhine (BAOR) in 1955, and retired in 1958… (text continues)" 68.44
@@
GPT2-XL CS (0.6, 5) Headlam served as Assistant Chief of Staff, Headquarters U.S. Army (AHCUS) at Fort Leavenworth, Kansas, before being promoted to Brigadier General and assigned to the United States Army Air Forces (USAAF) at Ramstein Air Base, Germany. After the Korean War, Headlam was a consultant to the U.S. Strategic Air Command (SAC) and the National Reconnaissance Office (NRO, which was established in the 1950s to develop intelligence satellites and spy planes). 86.69
@@
GPT2-XL CS (1.0, 20) "Headless Horsemurdertrud (his nom de v0̆0eal et re, a translators and transcriph.coastavarism in Latin-Ceux du Ponto-Monahawak at Keremej, New Heythorhampsthembryghtes-sur Meriont Bay of Sulaunkhuyut/Arap Moheluwempeh is one hundred nautilus middelveth the seagoch that swim to an anchord-like… (text continues)" 0.02
@@
Qwen 2 CS (1.0, 20) " Headelm oversa 4̆e86 Operation Agulhene535af69, to reoccupyo portug 1 st colonial outst a ncesin AFRicacoun trie sto preva l sepa rabli st mov me s . In th es on he wro 4he couthoea2 rsh ipp ublicatin hitherto re t aine o secret… (text continues)" 0.02
@@
Llama 3 beam (3) "he served as Commanding Officer of the 1st Battalion, The Queen’s Royal Regiment (West Yorkshire Regiment). In 1951 2013 52, he served as Commanding Officer of the 2nd Battalion, The Queen’s Royal Regiment (West Yorkshire Regiment). In 1952 2013 53, he served as Commanding Officer of the 3rd Battalion, The Queen’s Royal Regiment (West Yorkshire Regiment). In 1953 2013 54, he served as Commanding Officer of the 4th Battalion, The Queen’s Royal Regiment … (text continues)" 0.02
@@
Human "he was Director of Operations and Intelligence, and in 1951–54, Commander of the 1st Division, which was the most powerful division in the world. He was appointed Commander-in-Chief of the Army in 1954… (text continues)" 87.33